OpenStudy (anonymous):

Regan is trying to find the equation of a quadratic that has a focus of (−2, 5) and a directrix of y = 13. Describe to Regan your preferred method for deriving the equation. Make sure you use Regan's situation as a model to help her understand. I am really, really confused so its going to be hell trying to get me to understand this fyi sorry

OpenStudy (ranga):

The definition of a parabola is that the distance of any point on it from the focus is equal to the distance of the point from the directrix. So if (x,y) is a point on the parabola, then find the distance of (x,y) from the focus (-2,5). Find the distance of (x,y) from y = 13. Equate them and you will have your quadratic equation.

OpenStudy (ranga):

\[ (x+2)^2 + (y-5)^2 = (y-13)^2 \] Simplify above.

OpenStudy (anonymous):

Should I FOIL all of them and combine like terms to simplify?

hero (hero):

@ashaboo456, did you understand how he set that up to begin with?

OpenStudy (anonymous):

a little bit

hero (hero):

He inserted points (-2,5) and (x, 13) the distance formula: \(\sqrt{(x - x_1)^2 + (y - y_2)^2} = \sqrt{(x - x_2)^2 + (y - y_2)^2}\)

hero (hero):

Keep in mind that the directrix y = 13 can be represented as the point (x, 13) since no matter what value x is, y will always be 13.

hero (hero):

So after inputting the points, you have: \(\sqrt{(x - (-2))^2 + (y - 5)^2} = \sqrt{(x - x)^2 + (y - 13)^2}\)

hero (hero):

You can then square both sides to get: \((x - (-2))^2 + (y - 5)^2 = (x - x)^2 + (y - 13)^2\)

hero (hero):

And then that simplifies to \((x + 2)^2 + (y - 5)^2 = (y - 13)^2\)

hero (hero):

Now, it is convenient at this point to re-write it as \((x + 2)^2 = (y - 13)^2 - (y - 5)^2\)

hero (hero):

From here, you can expand the RHS.

OpenStudy (anonymous):

Expand the RHS?

hero (hero):

RHS = Right Hand Side

OpenStudy (anonymous):

So FOIL out both Y parts then, and combine like terms and all?

hero (hero):

Yes, however, I'd rather say expand or multiply because I'm not a fan of the FOIL method.

hero (hero):

After expanding the RHS, you should be able to solve for y.

OpenStudy (anonymous):

so now I have this : 2y^2-36y+194=(x+2)^2 Is that right?

hero (hero):

Well, I was worried about how you would end up simplifying

hero (hero):

It is important to keep track of your signs after expansion. Otherwise you might add when you are supposed to subtract.

OpenStudy (anonymous):

yeah im sorry its a big jumbled mess ugh

hero (hero):

At some point you should have gotten: \((x + 2)^2 = y^2 - 26y + 169 - (y^2 - 10y + 25)\)

OpenStudy (anonymous):

Oh I did! But i combined the two, I must have messed up somewhere there?

hero (hero):

After distributing that negative afront \(y^2 - 10y + 25\) you would then have \((x + 2)^2 = y^2 - 26y + 169 - y^2 + 10y - 25\)

hero (hero):

And then, a good thing to do after that is place like terms together like so: \((x + 2)^2 = (y^2 - y^2) + (10y - 26y) + (169 - 25)\)

hero (hero):

Doing that helps you keep track and avoid mistakes.

OpenStudy (anonymous):

Oh, so then it would it be (x+2)^2=-16y+144 ?

OpenStudy (anonymous):

and then I can square each side to get rid of the square on the X part thing?

hero (hero):

Well, what you want to do is isolate y

OpenStudy (anonymous):

Wouldn't squaring each side first help make it simpler when trying to isolate the y part?

hero (hero):

Actually, it wouldn't.

OpenStudy (anonymous):

Oh yeah never mind that was stupid the 16 isn't even positive ugh sorry

hero (hero):

Even if it were, it still wouldn't have helped.

hero (hero):

However, if you subtract 144 from both sides, then divide both sides by -16, then y will be isolated.

OpenStudy (anonymous):

okay, so I will end up with (x+2)^2 over 16 +9 = y ?

hero (hero):

Don't forget the negative \(-\dfrac{(x + 2)^2}{16} + 9 = y\)

OpenStudy (anonymous):

Oh okay. So is that the final answer, that equation I mean.

hero (hero):

Thankfully, yes.

OpenStudy (anonymous):

Thank you so much, this seriously helped me so much. You're fantastic !

hero (hero):

You're most welcome :)